# The Impact of Spectroscopic Redshift Errors on Cosmological Measurements

### v1

@@ Acknowledgments
SH, AR, JY and JPK, acknowledge support from the Swiss National Science Foundation (SNF) ‘Cosmology with 3D Maps of the Universe” research grant 200020_207379. This research was essentially supported through computational resources provided by the High Performance Computing (HPC) facility of University of Geneva (UNIGE), especially for the Bamboo cluster.
@@ Appendix A BAO fitting results
Catalogs clean dv-LRG dv-QSO-G dv-QSO-L dv-ELGcatas dv-slitless $\alpha_{\mathrm{iso}}$ $1.0031\pm 0.0045$ $1.0027\pm 0.0044$ $1.0026\pm 0.0045$ $1.0028\pm 0.0046$ $1.0034\pm 0.0046$ $1.0023\pm 0.0046$ $\alpha_{\mathrm{ap}}$ $1.001\pm 0.016$ $1.000\pm 0.015$ $0.996\pm 0.016$ $0.999\pm 0.015$ $1.000\pm 0.015$ $0.997\pm 0.016$ $\chi^{2}/\rm{dof}$ $4.51/33$ $5.45/33$ $6.25/33$ $5.89/33$ $4.67/33$ $5.23/33$
@@
Table 4: Comparison of BAO fitting results for the clean and contaminated mock catalogs described in Table˜1. Table reports the mean posterior values with 68% confidence intervals. The $\alpha_{\mathrm{iso}}$ and $\alpha_{\mathrm{ap}}$ remain almost unbiased in the presence of redshift errors ($\Delta\alpha_{\mathrm{iso}}<0.18\,\sigma$ and $\Delta\alpha_{\mathrm{ap}}<0.32\,\sigma$). The corresponding best-fit $\chi^{2}$ over the degree of freedom (dof) is also listed for each catalog, showing an increase when redshift errors are included.
@@
BAO serves as a standard ruler in cosmology and is generally considered robust against a wide range of systematic effects. In this work, we also test whether BAO measurements remain unbiased under the spectroscopic redshift errors. BAO analysis focus on constraining two dilation parameters, $\alpha_{\parallel}$ and $\alpha_{\perp}$, which are reparameterized into two scaling parameters, $\alpha_{\mathrm{iso}}$ and $\alpha_{\mathrm{ap}}$, respectively. These two parameters provide information about the isotropic BAO scale and AP effect. We adopt the BAO modeling framework implemented in DESI DR1 analyses [68] using the desilike package. The results are obtained by fitting the mean monopole and quadrupole of the two-point correlation function from 500 mock realizations without BAO reconstruction. The built covariance matrix are rescaled to an effective volume of $V_{25}=25\,h^{-3}\rm{Gpc}^{3}$.
Table˜4 summarizes the resulting BAO constraints. We find that the differences in $\alpha_{\mathrm{iso}}$ and $\alpha_{\mathrm{ap}}$ estimates remain small, with $\Delta\alpha_{\mathrm{iso}}<0.18\,\sigma$ and $\Delta\alpha_{\mathrm{ap}}<0.32\,\sigma$, and the corresponding error bars remain nearly unchanged. The best-fit $\chi^{2}$ values increase noticeably after including redshift errors, with up to a 40% increase observed in the dv-QSO-L sample. This indicates a decrease of the goodness of fit due to the redshift errors. But in general, the BAO measurements remain unbiased and robust against spectroscopic redshift errors at the current level of precision.
Appendix B $\chi^{2}$ results and contour plot for baseline analysis
@@
Figure 9: Comparison of corresponding $\chi^{2}$ for Figure˜4 and Figure˜5. The best-fit $\chi^{2}$ values are normalized by the degrees of freedom, with results for ShapeFit (left panels) and Full Modeling (right panels). For each point, we fit the mean of 25 random mocks, with the covariance rescaled to $V_{25}$. The difference of $\chi^{2}$ is small between clean and contaminated mock catalogs, demonstrating that all fits achieve the same level of goodness.
Figure 10: Posterior distribution for nuissance parameters in our baseline EFT  fitting. Contours show the 68% and 95% confidence interval. Results are derived from fits to the mean power spectrum monopole and quadrupole of random selected 25 mocks using the covariance matrix rescaled to $V_{25}$. The negative shift of $\alpha_{2}$, $\alpha_{4}$ on their marginalized posterior are account for the effect of redshift uncertainty, especially for the QSO-like smearing cases.
@@
Figure˜9 presents the distribution of best-fit $\chi^{2}$ values for the baseline analysis. The left two panels show results for the ShapeFit analysis, corresponding to the scatter plot in Figure˜4, while the right two panels show the Full-Modeling results for Figure˜5. From the $\chi^{2}$ distributions, we do not observe significant differences ($15\%$) or systematic shifts between the clean and contaminated mocks analyses. This indicates that the theoretical model provides a same level of good prediction with the observational data, even potential deviations in the parameter estimates may still arise under contamination.
Figure˜10 shows a representative posterior distribution for nuisance parameters in EFT of LSS framework, comparing results from the clean mocks and those contaminated by redshift uncertainty. We recall that the $b_{i}$ parameters correspond to the biased terms, $\alpha_{i}$ denote the EFT counterterms and $s_{n,i}$ model the stochastic contribution. The most noticeable shifts in the posteriors are observed for $\alpha_{2}$, $\alpha_{4}$, and $s_{n,2}$, especially for the QSO-like smearing cases. As explained in Section˜4.3.1, the EFT counterterms $\alpha_{2}$ and $\alpha_{4}$ are expected to absorb the damping effect in the power spectrum introduced by redshift uncertainty. This behavior is reflected as a negative shift in their estimated values, with the value increasing with the level of smearing. The parameter $s_{n,2}$ represents the second-order stochastic term, which scales as $(k\mu)^{2}$. Its variations likely reflect the stochastic fluctuations during the fitting, but we do not take them explicitly accounted for redshift uncertainty effect.

### v4

@@ Acknowledgments
SH, AR, JY and JPK, acknowledge support from the Swiss National Science Foundation (SNF) ‘Cosmology with 3D Maps of the Universe” research grant 200020_207379. This research was essentially supported by computational resources provided by the High Performance Computing (HPC) facility of University of Geneva (UNIGE), especially for the Bamboo cluster.
@@ Appendix A BAO fitting results
Catalogs clean dv-LRG dv-QSO-G dv-QSO-L dv-ELGcatas dv-slitless $\alpha_{\mathrm{iso}}$ $1.003\pm 0.005$ $1.003\pm 0.005$ $1.002\pm 0.005$ $1.003\pm 0.004$ $1.003\pm 0.005$ $1.003\pm 0.005$ $\alpha_{\mathrm{ap}}$ $1.002\pm 0.016$ $0.998\pm 0.016$ $0.998\pm 0.016$ $0.999\pm 0.017$ $1.000\pm 0.017$ $0.998\pm 0.017$ $\chi^{2}/\rm{dof}$ $1.102\pm 0.274$ $090\pm 0.250 1.1$$109\pm 0.259$ $1.064\pm 0.229$ $1.074\pm 0.222$ $1.061\pm 0.230$
@@
Table 4: Comparison of BAO best-fit results for the clean and contaminated mock catalogs described in Table˜1. Table reports the mean and standard deviation of best-fit values from 200 fits. The $\alpha_{\mathrm{iso}}$ and $\alpha_{\mathrm{ap}}$ constraints remain unbiased in the presence of redshift errors. The corresponding best-fit $\chi^{2}$ over the degree of freedom (dof) shows the same level of goodness-of-fit.
@@
BAO serves as a standard ruler in cosmology and is generally considered robust against a wide range of systematic effects. In this work, we also test whether BAO measurements remain unbiased under the spectroscopic redshift errors. BAO analysis focuses on constraining two dilation parameters, $\alpha_{\parallel}$ and $\alpha_{\perp}$, which are reparameterized into two scaling parameters, $\alpha_{\mathrm{iso}}$ and $\alpha_{\mathrm{ap}}$, respectively. These two parameters provide information about the isotropic BAO scale and AP effect. We adopt the BAO modeling framework implemented in DESI DR1 analysis [desi_collaboration_desi_2025-2] using the desilike package. The results are obtained by fitting the monopole and quadrupole of the two-point correlation function from the mean of 25 randomly selected mocks out of 500 realizations without BAO reconstruction. The covariance matrix is built from 500 mocks and rescaled to an effective volume of $V_{25}=25\,h^{-3}\rm{Gpc}^{3}$.
Table˜4 summarizes the mean and standard deviation for $\alpha_{\mathrm{iso}}$ and $\alpha_{\mathrm{ap}}$ best-fit values and $\chi^{2}$ obtained from 200 independent fits. The inferred $\alpha_{\mathrm{iso}}$ and $\alpha_{\mathrm{ap}}$ values show only minor differences from the clean case, both in their means and standard deviations. The best-fit $\chi^{2}$ values also show no increase after including redshift errors, indicating a same level of goodness of fit. Therefore, we conclude that BAO measurements remain unbiased and robust against spectroscopic redshift errors at the current level of precision.
Appendix B $\chi^{2}$ results for baseline analysis
@@
Figure 10: Comparison of corresponding $\chi^{2}$ for Figure˜5 and Figure˜6. The best-fit $\chi^{2}$ values are normalized by the degrees of freedom, with results for ShapeFit (left panels) and Full Modeling (right panels). For each point, we fit the mean of 25 random mocks, with the covariance rescaled to $V_{25}$. The difference of $\chi^{2}$ is small between clean and contaminated mock catalogs, demonstrating that all fits achieve the same level of goodness.
@@
Figure˜10 presents the distribution of best-fit $\chi^{2}$ values for the baseline analysis. The left two panels show results for the ShapeFit analysis, corresponding to the scatter plot in Figure˜5, while the right two panels show the Full-Modeling results for Figure˜6. From the $\chi^{2}$ distributions, we do not observe significant differences ($15\%$) or systematic shifts between the clean and contaminated mocks analyses. This indicates that the theoretical model provides a same level of good prediction with the observational data, even potential deviations in the parameter estimates may arise under contamination.